# AI-driven Wireless Positioning: Fundamentals, Standards, State-of-the-art, and Challenges

### v1

@@ VIII-A1 Data Collection
The challenges in data collection for AI-driven wireless positioning are multifaceted, stemming from the dynamic, diverse, and sensitive nature of wireless data. High-quality labeled datasets are essential but difficult to obtain due to the diverse propagation characteristics across environments (urban, rural, indoor, outdoor) and the dynamic nature of wireless signals [361]. Datasets quickly become outdated, necessitating frequent updates, while over-reliance on synthetic datasets often fails to capture real-world complexities, leading to reduced model robustness. Furthermore, the discrepancies between simulated and real-world channel characteristics make it difficult for models trained on synthetic data to perform effectively in practical environments, significantly increasing the workload associated with real-world data collection. Additionally, the massive volume and high dimensionality of wireless data, particularly from advanced technologies like massive MIMO and mmWave, pose computational challenges. Outliers and noise in raw data further degrade model performance, requiring robust preprocessing and anomaly detection techniques [362].
@@ VIII-A2 Accuracy in Complex Environments
Achieving high positioning accuracy in complex environments remains a significant challenge for AI-driven wireless positioning systems. Real-world scenarios such as urban areas, highways, and indoor settings are characterized by dynamic and unpredictable factors that reduce the precision of positioning algorithms. These factors include multipath propagation, NLOS conditions, high-speed mobility, interference, and environmental obstacles. Under these circumstances, signals are reflected, refracted, and scattered, leading to complex and distorted channel distributions. While AI models are capable of learning intricate patterns, they require extensive and diverse data to generalize effectively in such environments. However, collecting such data is often impractical or infeasible. Consequently, the challenge of achieving reliable accuracy in complex and dynamic environments continues to impede the practical deployment of AI-driven wireless positioning systems.
@@ VIII-A3 Model Generalization
As with most AI applications, one of the key challenges in AI-driven wireless positioning is ensuring model generalization across diverse environments and deployment scenarios. Variations in signal propagation characteristics, device configurations, and environmental dynamics often cause AI models trained under specific conditions to perform poorly when applied to new or unseen environments. For instance, a model trained in an urban setting may struggle in rural or indoor environments where channel characteristics differ significantly. While advanced techniques such as domain adaptation, transfer learning, and multi-task learning can partially address this issue, their ability to fully resolve the complexity and dynamic nature of wireless positioning environments remains limited.
@@ VIII-A4 Resource Constraints
Resource constraints pose a significant barrier to the deployment of AI-based positioning systems, particularly on edge devices such as smartphones, IoT sensors, and UAVs. AI model inference and training typically demand substantial computational power, memory, and energy, which often exceed the capabilities of these devices. Additionally, applications requiring real-time positioning, such as autonomous driving, face added complexity when processing high-dimensional data from advanced technologies like massive MIMO and mmWave systems. Developing lightweight models and efficient algorithms that balance accuracy and computational efficiency is critical but remains a challenging endeavor.
VIII-A5 Integration with Emerging Wireless Technologies
The integration of AI-driven positioning systems with emerging wireless technologies such as RIS, massive MIMO, THz communications, and NTN introduces significant challenges. These technologies offer unique capabilities, such as enhanced spatial resolution and broad coverage, which can greatly improve positioning accuracy. However, leveraging these advantages requires the development of new AI models and algorithms, adding complexity to their seamless integration with positioning systems. Additionally, designing scalable AI models capable of adapting to the diverse characteristics of different wireless technologies remains a substantial challenge, further complicating their deployment in real-world scenarios.
@@ VIII-A6 Security and Privacy Concerns
Security and privacy are critical concerns in AI-driven wireless positioning. Raw data used for positioning, such as CSI, often contains sensitive user information, including precise locations and device identifiers, making it vulnerable to misuse, tampering, or leakage during storage and transmission. Positioning systems are also susceptible to attacks like spoofing, interference, and eavesdropping, which can compromise their reliability and accuracy. Additionally, the privacy of AI models themselves must be safeguarded. Privacy-preserving techniques such as federated learning and differential privacy offer promising solutions, but their deployment introduces additional computational overhead and complexity. Striking a balance between security and privacy requirements and maintaining system performance remains an ongoing challenge.
@@ VIII-B Opportunities in AI-Driven Wireless Positioning
While wireless AI positioning faces numerous challenges, it also presents unprecedented opportunities for innovation and application across various domains. These opportunities arise from advancements in AI technologies, the evolution of wireless networks, and the growing demand for high-precision positioning solutions.
@@ VIII-B1 Enhanced Accuracy Through Advanced AI Technologies
In recent years, AI technologies have advanced rapidly, presenting immense potential for overcoming the limitations of traditional wireless positioning. Cutting-edge AI models and techniques, such as Transformers, self-supervised learning, and large language models (LLMs), are continuously emerging. These innovations enable AI systems to model the nonlinear dynamics of signal propagation, achieving centimeter-level precision in challenging scenarios. For instance, Transformers can be employed to capture spatiotemporal features of wireless channels in the time and frequency domains, while graph neural networks enhance model scalability and adaptability. Moreover, self-supervised learning methods allow AI systems to extract latent representations from unlabeled datasets, creating robust models even in the absence of extensive labeled data. With the advancement of AI technology, advanced AI algorithms will have great potential to gradually solve various problems encountered in wireless positioning.
VIII-B2 Multimodal Data Fusion for Robust Systems
Advancements in LLMs have unlocked transformative opportunities for AI-driven wireless positioning by enabling the fusion of multiple data modalities. By integrating wireless signal data with additional inputs such as visual data, inertial measurements, and environmental context, AI-driven systems can achieve exceptional robustness and adaptability. For instance, combining CSI with vision-based inputs can significantly improve positioning accuracy in indoor environments. Furthermore, the fusion of LLM technology with wireless signal data introduces novel capabilities for wireless networks, such as enabling natural language-based navigation and query processing applications. This multimodal approach not only enhances system resilience in complex scenarios but also expands the application scope of AI-driven wireless positioning systems.
VIII-B3 Advances in Emerging Wireless Technologies
Emerging wireless technologies such as massive MIMO, RIS, NTN, and THz communications open up new research directions for AI-driven positioning. These technologies introduce novel features to wireless networks that create exciting opportunities for enhanced localization capabilities. For example, RIS offers programmable propagation environments, enabling dynamic control over signal paths, while massive MIMO arrays provide exceptionally high spatial resolution. These advancements complement AI algorithms, allowing them to leverage these unique characteristics to deliver superior positioning accuracy.
VIII-B4 Resource-Efficient AI Models
@@
The development of lightweight and resource-efficient AI models offers a promising pathway for the widespread deployment of wireless positioning systems on resource-constrained devices. Techniques such as model pruning, quantization, and knowledge distillation enable the creation of compact AI models that reduce computational and energy requirements while maintaining high accuracy. Achieving precise positioning with small-scale models remains a crucial and impactful research challenge. Addressing this can facilitate the integration of AI-driven positioning systems into IoT devices, wearable technologies, and mobile platforms.
@@ VIII-B5 ISAC-Enhanced AI Positioning
integrated sensing and communication (ISAC) systems [363] enhance positioning by utilizing echo signals to sense the surrounding environment and extract key parameters such as AoA, ToA, and Doppler shift. This capability provides a deeper understanding of the network environment, such as NLOS detection [364]. By incorporating sensing signals alongside communication-based positioning methods, ISAC systems can significantly enhance the accuracy and robustness of wireless positioning. In such a complex system, AI algorithms play a pivotal role by leveraging their advanced feature extraction capabilities to process the multimodal data from ISAC systems, enabling high-precision positioning even in challenging scenarios.

### v2

@@ VII-A7 ViWi
VII-B Case Study in AI-Driven Positioning
Figure 10: Localization performance comparison using the MaMIMO dataset under ULA-LOS configuration.
Figure 11: Localization performance comparison using the DeepMIMO ‘O1’ scenario.
To evaluate the impact of different datasets and models on positioning performance, we conduct a comparative case study using both the measured MaMIMO dataset and the simulated DeepMIMO dataset. For the MaMIMO dataset, we focus on the ULA configuration under LOS conditions. For DeepMIMO, we select the widely used urban street scenario "O1", which is a typical deployment environment. In both datasets, we randomly sample 7,000 samples for training, 1,000 for validation, and 2,000 for testing. To ensure a fair comparison, we fix key channel parameters across both scenarios: 20 MHz bandwidth, 100 OFDM subcarriers, and a 64-element ULA at the BS. We evaluate four neural network backbones: FCNN, CNN (ResNet-32), LSTM, and Transformer. Each model is trained and tested under the same data split for both datasets.
The results are summarized in Fig. 10 and Fig. 11. The models exhibit different performance across the two datasets. In the MaMIMO scenario, the denser and more concentrated spatial sampling allows the models to achieve higher positioning accuracy. Nevertheless, the ranking of model performance is consistent across both datasets: from best to worst, the order is CNN, Transformer, LSTM, and FCNN. Although the Transformer demonstrates stronger feature representation capabilities, it requires more data for training to fully converge. Therefore, it performs worse than CNN in situations with limited data or without pre-training. Specifically, in the MaMIMO dataset, the median positioning errors achieved by FCNN, CNN, LSTM, and Transformer are 141.2 mm, 33.1 mm, 51.8 mm, and 40.9 mm, respectively. In the DeepMIMO dataset, the corresponding median errors are 13.1 m, 1.0 m, 1.7 mm, and 1.2 m, respectively.
VII-C Future Directions for Dataset Design
Despite the growing availability of wireless positioning datasets, there remains no unified benchmark that supports dynamic, multi-modal, and heterogeneous conditions. Most existing datasets suffer from limitations such as lack of user mobility, restricted environmental diversity, or absence of cross-modal sensing data, which hinders the generalizability and reproducibility of AI-based localization solutions. Therefore, future dataset development should consider the following directions:
  • Multi-modality integration: Combine RF data with auxiliary sensing modalities such as IMU, LiDAR, cameras, and digital maps to enable multi-sensor fusion, learning-based SLAM, and cross-domain localization tasks.
  • Doppler and time-evolving channels: Incorporate realistic time-series CSI that captures Doppler effects, user mobility, and handover dynamics to support mobility-aware learning and trajectory prediction.
  • Scalable and customizable simulation: Develop open-source, extensible simulation platforms (e.g., built on Unity or Unreal Engine) that allow researchers to design custom virtual environments and generate synchronized RF and sensor data in real time.
  • Alignment between synthetic and real-world data: Promote methodologies for bridging the gap between simulated datasets and real-world environments, enabling models trained in virtual domains to generalize effectively to physical deployments.
  • Standardized benchmarks and evaluation metrics: Establish public benchmarks with consistent data splits, task definitions, and evaluation metrics (e.g., accuracy, latency, energy efficiency) to foster reproducibility and fair comparison across algorithms.
VII-D Lessons Learned
We reviewed representative datasets, including MaMIMO, DeepMIMO, WAIR-D, and xG-Loc, covering both measured and simulated data sources. Dataset design significantly influences the performance of AI-based positioning models. Measured datasets like MaMIMO offer higher positioning accuracy due to dense and precise spatial sampling, while simulated datasets such as DeepMIMO and WAIR-D provide configurability but lack temporal channel dynamics and real-world variability. Moreover, few existing datasets support user mobility, multi-band signals, or cross-modal data (e.g., vision or IMU), limiting their applicability in complex 6G scenarios. Our case studies further show that CNNs perform consistently well on DeepMIMO and MaMIMO datasets, while Transformers require larger datasets to achieve optimal performance. Overall, current datasets fall short in enabling generalizable, dynamic, and multi-modal learning, underscoring the urgent need for standardized and extensible datasets to benchmark and accelerate wireless positioning research for future networks.
@@ VIII-A1 Data Collection
Data collection for AI-driven wireless positioning presents unique challenges, primarily due to the scenario-specific nature of positioning tasks, which necessitates customized datasets. This significantly increases the difficulty of acquiring accurate ground-truth position labels. Heterogeneous propagation characteristics across different environments (e.g., urban canyons, dense indoor layouts, rural open spaces) often lead to severe performance degradation when a model trained in one setting is transferred to another. Additionally, the dynamic nature of wireless propagation causes fingerprints to become outdated quickly, thereby increasing the demand for frequent dataset updates and incurring substantial maintenance costs.
High-quality labeled datasets are essential for supervised learning models. To overcome this challenge, one promising direction is the development of digital twin systems tailored for wireless positioning [202]. These systems leverage ray tracing techniques to generate synthetic wireless channels that can be used for model training. However, a major limitation lies in the domain gap between simulated and real-world channels, which undermines model generalization and highlights the need for further advancement in accurate channel modeling. An alternative approach involves the use of generative neural networks. GANs have already demonstrated strong performance in existing wireless positioning works [143]. As AI technology continues to evolve, more advanced generative models, such as diffusion models and large AI models, are expected to further bridge the gap between synthetic and real-world data. These techniques provide a promising solution for building data-efficient and scalable wireless positioning systems that work well in diverse and dynamic environments.
@@ VIII-A2 Accuracy in Complex Environments
Achieving high positioning accuracy in complex and high-interference environments remains a critical challenge for AI-driven wireless positioning systems. Real-world scenarios such as urban canyons, highways, dense indoor areas, or industrial facilities are characterized by dynamic and unpredictable signal conditions, including multipath propagation, NLOS effects, high-speed mobility, and external interference. In such environments, wireless signals are frequently reflected, refracted, or scattered, resulting in distorted and non-stationary channel characteristics. These factors not only degrade signal quality but also lead to rapid shifts in the underlying data distribution, making it difficult for AI models. Moreover, the presence of strong interference from overlapping wireless systems or co-channel users further reduces signal fidelity, introducing noise that hampers both model training and inference. While AI models have demonstrated the capability to learn complex propagation patterns, they often require large volumes of diverse and up-to-date labeled data to generalize effectively. However, such data is costly and often infeasible to collect at scale in dynamically evolving environments.
To address this, one promising direction is to leverage foundation models pretrained on large-scale unlabeled wireless data, which can enhance the representation capability of AI models and improve localization accuracy [201]. Researchers are also exploring hybrid approaches that combine data-driven learning with model-based methods, aiming to improve robustness in complex environments [331]. Moreover, digital twin systems can simulate intricate propagation conditions ahead of deployment, enabling the design of data- and model-driven positioning algorithms that are more resilient and adaptable to real-world scenarios [202]. Finally, multi-modal positioning [359], which incorporates complementary sensory data such as inertial measurements, camera input, or geomagnetic signals, can help compensate for radio signal degradation, thus improving accuracy and reliability in harsh environments.
@@ VIII-A3 Model Generalization
Model generalization remains one of the most critical challenges in AI-driven wireless positioning, particularly due to the high variability of deployment environments and signal propagation conditions. Wireless environments vary significantly across different geographic locations, building layouts, materials, hardware platforms, and time periods. For example, models trained in urban environments may fail to perform in rural areas or indoor industrial settings due to stark differences in channel characteristics, such as multipath density and obstruction patterns. These discrepancies result in domain shifts in both feature space (e.g., CSI or RSSI distributions) and label space (e.g., spatial layouts), making it difficult for AI models to maintain robust performance across unseen scenarios.
While advanced techniques such as domain adaptation, transfer learning, and multi-task learning can partially address this issue, their ability to fully resolve the complexity and dynamic nature of wireless positioning environments remains limited. To further enhance model generalization, one promising approach is to leverage federated learning frameworks that aggregate heterogeneous channel data from diverse deployment scenarios, enabling collaborative training [360]. Additionally, large AI foundation models pretrained on large-scale unlabeled wireless datasets have the potential to learn domain-invariant and generalizable positioning representations, thereby improving adaptability to unseen environments [201]. Finally, model-based approaches that incorporate digital twin systems and physical prior knowledge, such as floor plans [336] or environmental maps [338], can also help bridge the generalization gap by introducing environment information into the learning process.
@@ VIII-A4 Resource Constraints
Resource constraints pose a significant barrier to the deployment of AI-based positioning systems, particularly on edge devices such as smartphones, IoT sensors, and UAVs. AI model inference and training typically demand substantial computational power, memory, and energy, which often exceed the capabilities of these devices. Additionally, processing wireless positioning data (e.g., high-dimensional CSI matrices from massive MIMO or raw CIR signals from mmWave systems) typically involves deep neural networks with millions of parameters, resulting in high inference latency and power consumption. This challenge is particularly severe in latency-critical applications like autonomous vehicles or AR/VR. Moreover, model updates in dynamic environments demand continual learning or on-device adaptation, further straining computational budgets.
To address these issues, model compression techniques such as weight pruning [361], quantization [362], and knowledge distillation [363] can significantly reduce model size and computation cost while maintaining accuracy. Beyond compression, Edge-AI paradigms have demonstrated great potential. In particular, federated learning [364] enables collaborative model training across distributed devices without transmitting raw data, thereby reducing communication costs and preserving user privacy. Meanwhile, distributed inference techniques [365, 203] leverage the computing power of nearby edge servers to offload heavy computation, alleviating the burden on individual devices. These approaches together offer a promising direction for enabling scalable and efficient AI-driven wireless positioning in constrained environments.
VIII-A5 Scalability Across Network Configurations
Scalability across different wireless network configurations presents another key challenge for AI-driven positioning systems. On one hand, scalability across BS configurations refers to the ability of AI models to operate effectively under diverse network settings. In practice, BSs may employ different configurations in terms of carrier frequency (e.g., sub-6 GHz vs. mmWave), bandwidth, MIMO setup, antenna layout, and pilot signaling schemes to meet varying deployment requirements. These differences lead to significant variability in the format and characteristics of the CSI, such as input dimensionality, spectral properties, and temporal structure. As a result, models trained under one configuration often struggle to generalize across others, reducing their transferability and deployment flexibility. On the other hand, scalability at the system level involves handling large-scale, complex network scenarios, such as multi-BS, multi-user interference, and ultra-dense infrastructure deployments. Emerging technologies like massive MIMO, RIS, and NTN networks introduce more intricate propagation behaviors and generate vast amounts of high-dimensional data. These challenges highlight the necessity for flexibly scalable AI algorithms that can adapt to heterogeneous network parameters and system conditions.
To address the scalability challenge, future solutions may leverage large AI foundation models tailored for positioning tasks. By pretraining on large-scale unlabeled CSI data from diverse network configurations, these models can learn domain-invariant representations that align features across different BS settings and infrastructures (e.g., RIS, NTN). This enables the model to generalize across varying input formats and supports multi-source CSI fusion, improving adaptability and robustness in complex deployment scenarios.
@@ VIII-A6 Security and Privacy Concerns
Security and privacy are also very important for positioning applications in future networks [5]. AI-based wireless positioning systems face three primary categories of security and privacy risks: (1) Vulnerabilities during signal acquisition: Raw positioning signals, such as CSI and RSS, can be intercepted, spoofed, jammed, or perturbed by adversaries. These threats degrade signal integrity and can significantly undermine localization accuracy, especially in dynamic or hostile environments [366, 367]. (2) Privacy leakage during model training: The training phase often involves large volumes of sensitive user data containing precise location traces. Such data is vulnerable to privacy attacks, including data reconstruction, membership inference, and gradient leakage, which can reveal personal information or re-identify individuals from aggregated datasets. (3) Attacks during model inference: The AI models themselves may be exploited as attack vectors. Techniques such as model inversion, membership inference, or backdoor injection can be used to extract private information or manipulate the model’s output. In addition, attackers may introduce imperceptible perturbations to input signals, tricking the model into outputting incorrect or misleading positions.
To mitigate the security and privacy risks in wireless positioning, several countermeasures can be applied at different stages of the system [5]. During the signal acquisition phase, physical-layer security techniques can be employed to detect and exclude adversarial sources, thereby mitigating spoofing and jamming attacks. In addition, signal preprocessing and robust feature extraction methods can reduce the susceptibility of raw signals to malicious perturbations [368]. During the model training phase, federated learning offers a promising solution by enabling decentralized model updates without sharing raw location data, thus preserving user privacy while also improving model generalization [364]. For threats occurring during inference, techniques such as differential privacy can increase the difficulty of model inversion, membership inference, or backdoor injection attacks. Furthermore, incorporating adversarial training during the learning process can enhance model robustness against input perturbations and adversarial manipulation [367].
VIII-A7 Explainability and Trustworthiness
AI-driven wireless positioning systems often rely on complex deep learning models that function as black boxes with limited interpretability. This lack of transparency presents a major obstacle to practical deployment, particularly in safety-critical applications such as autonomous navigation, emergency response, and industrial automation, where black-box AI algorithms cannot provide sufficient guarantees of reliability. Without clear explanations, it becomes difficult to trace, debug, or correct erroneous outputs, thereby undermining user confidence and accountability [369]. Therefore, understanding the underlying rationale behind model predictions is essential for building trustworthy systems.
To address this challenge, one direction is to pursue rigorous theoretical analysis of AI model reliability, although this remains one of the most difficult open problems in the research community. On the other hand, a more practical approach is to develop data-model hybrid AI algorithms that combine data-driven learning with physics-based models. By integrating domain knowledge and physical priors [4], these hybrid methods can enhance interpretability and improve the trustworthiness of AI-driven positioning systems.
@@ VIII-B Opportunities in AI-Driven Wireless Positioning
While wireless AI positioning faces numerous challenges, it also presents unprecedented opportunities for innovation and application across various domains. These opportunities arise from advancements in AI technologies, the evolution of wireless networks, and the growing demand for high-precision positioning solutions. Building on the aforementioned challenges and their potential solutions, we outline below several key opportunities emerging in this field:
@@ VIII-B1 Enhanced Accuracy Through Advanced AI Technologies
Recent progress in AI has opened new possibilities for improving the accuracy of wireless positioning systems. Advanced models such as Transformers, self-supervised learning, and large-scale foundation models have shown strong capabilities in capturing the nonlinear and complex nature of wireless signal propagation. Especially, inspired by the success of foundation models in vision and language, similar paradigms are now emerging in wireless communications. By pretraining on large-scale unlabeled datasets from diverse environments, these models can learn transferable and generalizable channel semantics, significantly reducing reliance on labeled data and improving robustness across deployment scenarios [201]. Moreover, their scalability enables them to handle heterogeneous inputs, including varying frequency bands, antenna configurations, signaling protocols, and even emerging technologies such as RIS and NTN. They can also integrate multi-modal data (e.g., vision and LiDAR), making them highly adaptable to diverse and evolving wireless infrastructures.
VIII-B2 Multi-Source and Cross-Modal Data Fusion
AI-driven wireless positioning is increasingly benefiting from the fusion of data collected across multiple network infrastructures and sensing modalities. Emerging network architectures, such as cell-free massive MIMO, NTN, and UAV-assisted networks, offer complementary spatial and temporal coverage, enriching the available wireless signal space for localization tasks. Simultaneously, multi-modal sensor data, including vision, LiDAR, and inertial measurements, provides additional context that can compensate for the limitations of radio-based methods under challenging conditions such as NLOS or deep indoor environments. By leveraging AI models capable of cross-modal representation learning and feature alignment, positioning systems can integrate heterogeneous data sources to achieve more robust, accurate, and environment-aware localization. Moreover, large AI foundation models provide a unified framework for learning consistent representations across diverse modalities and input formats. Their capacity for multi-source feature abstraction and semantic alignment enables deeper fusion at both the feature and decision levels, laying the foundation for scalable and generalizable positioning systems in complex, multi-modal environments.
VIII-B3 Data-Model Co-Design through Digital Twins
The emergence of digital twin technologies introduces a promising paradigm for both dataset generation and the co-design of data- and model-driven wireless positioning algorithms [202, 370]. By simulating realistic radio environments through ray tracing and physical channel modeling, digital twins enable the creation of large-scale, scenario-specific synthetic datasets, significantly reducing the cost and effort of data collection. These virtual environments also support controlled model pretraining, rigorous evaluation under diverse conditions, and reproducible benchmarking. Beyond data generation, digital twins serve as a bridge between physics-based modeling and AI learning frameworks. By embedding physical priors (such as geometry, material properties, and propagation constraints) into AI models [338], they enable hybrid learning strategies that fuse model knowledge with data-driven optimization. This co-design approach enhances model interpretability, generalizability, and robustness.
VIII-B4 Resource-Efficient AI Models and Edge Deployment
@@
The development of lightweight and resource-efficient AI models offers a promising pathway for the widespread deployment of wireless positioning systems on resource-constrained devices. Techniques such as model pruning, quantization, and knowledge distillation enable the creation of compact AI models that reduce computational and energy requirements while maintaining high accuracy. In addition, offloading positioning tasks to edge servers or adopting distributed inference frameworks allows for efficient utilization of edge computing resources, alleviating the processing burden on local devices. As a result, resource-efficient AI and edge deployment together enable the practical integration of AI-driven positioning into a broader range of applications, even on devices with limited hardware capabilities such as IoT nodes, wearable devices, and mobile platforms.
@@ VIII-B5 ISAC-Enhanced AI Positioning
Integrated sensing and communication (ISAC) systems [371, 372] enhance positioning by utilizing echo signals to sense the surrounding environment and extract key parameters such as AOA, TOA, and Doppler shift. This capability provides a deeper understanding of the network environment, such as NLOS detection [373]. ISAC systems can infer the position and motion of surrounding objects even without explicit transmissions from them. In complex and dynamic scenarios, AI can be used to jointly extract and fuse spatial features from both positioning and sensing modalities, enabling high-precision localization through the synergy of active and passive information. This opens up new possibilities for robust, environment-aware, and context-adaptive positioning in future wireless networks.
VIII-B6 Privacy- and Security-Aware AI Positioning
As AI-driven wireless positioning systems are increasingly deployed in user-centric and safety-critical scenarios, ensuring data privacy and system security has become a fundamental design requirement. This opens new opportunities for building AI frameworks that are both privacy-preserving and resilient to attacks. It is essential to address not only the inherent privacy and security challenges of the positioning task itself, but also the vulnerabilities introduced by AI models during training and inference. Techniques like federated learning [374] and differential privacy enable collaborative model training without transmitting raw location data, significantly reducing the risk of data leakage. Meanwhile, adversarial training and robust feature extraction methods can enhance model resistance to spoofing, jamming, and model inversion attacks. By jointly considering the threats from both wireless data and AI components, future systems can achieve secure and trustworthy AI-based positioning.
VIII-B7 Robust and Trustworthy AI for Mission-Critical Applications
AI-based wireless positioning is becoming increasingly relevant in safety-critical applications such as emergency response, industrial automation, and autonomous driving. This trend underscores both the need and the opportunity to develop AI models that are not only accurate but also trustworthy and interpretable. Research on robust AI, adversarial defenses, and explainable machine learning can help build reliable systems capable of operating under uncertain or adversarial conditions. Furthermore, the integration of data- and model-driven algorithm design offers a promising path toward enhancing both the transparency and reliability of AI-based localization